# Automated Statistical Model Discovery with Language Models

### v1

@@ 3.1 Searching over a DSL: automated Gaussian process kernel discovery
Dataset LM+ LM Periodic AS SM N-BEATS
@@
Figure 2: Test set performance of LM system on time series datasets. The LM system identifies compositional kernels that match the performance of strong baselines, such as the Automated Statistician. (left) Comparison of Test Mean Absolute Error (MAE) against Automated Statistician (Greedy Search), Spectral Mixture kernel, Periodic kernel, and N-BEATS. LM+ refers to search over augmented base kernel space. We bold the best and underline second among GP methods. (right) Extrapolations from GP with LM discovered kernel.
We first evaluate LM’s ability to search over a constrained space of models; in some settings, a domain expert may require the modeler to search over a DSL for reasons such as interpretability. In particular, we consider time-series modeling with Gaussian processes (GPs) as in the Automated Statistician (Duvenaud et al., 2013). GPs are probabilistic models that specify a distribution over functions. GPs are defined by a mean function and a positive-definite kernel function $k(\mathbf{x},\ \mathbf{x^{\prime}})$ that gives the covariance between $f(\mathbf{x})$ and $f(\mathbf{x^{\prime}})$ as a function of $\mathbf{x}$ and $\mathbf{x^{\prime}}$. The properties of a GP depend significantly on the kernel and therefore choosing a kernel that reflects domain knowledge (e.g., linearity, periodicity) is a key design choice. One common way to produce a more flexible kernel is to compose kernels via addition and multiplication, leveraging the closure properties of kernels. Here, we evaluate LMs ability to search over a space of kernels to identify an appropriate composition of kernels.
@@
Mirroring Duvenaud et al. (2013), we define a space of base kernels. In the prompt, we ask LM to perform one of the three operations (addition, multiplication, and replacement) on one of the in-context exemplar programs. For the first round, we ask the LM to produce an initial guess based on the structure of the dataset, which we provide as a plot. Given a kernel expression, we learn the kernel hyperparameters via gradient-based optimization of the marginal likelihood.
@@ Results
We evaluate our method on six common univariate time series datasets. We compare against the Automated Statistician, a greedy algorithm proposed by Duvenaud et al. (2013) (run until a depth of 10), and two GP baselines: a periodic kernel, a spectral mixture kernel (a strong non-compositional baseline). We also compare against N-BEATS, a strong neural baseline (Wilson & Adams, 2013; Oreshkin et al., 2019). In Figure 2, we compare the mean absolute error (MAE) on held-out test data for all datasets. To account for stochasticity in the Automated Statistician implementation we used (Saad et al., 2023) and stochasticity in different repetitions of our pipeline, we average the test MAE across three model runs. Our method, denoted LM in the table, matches the performance of the Automated Statistician, showing that the LM can efficiently search over a constrained space of models. We also experiment with augmenting the base kernel space with additional kernels (denoted LM+ in the table). In some cases, the LM can leverage this additional flexibility; for example, by using the additional kernels, LM+ can better capture the Australian beer sales dataset than LM. On the right panel, we show the extrapolations and identified kernel for the monthly air passenger dataset; LM+ identifies a kernel with a periodic times linear component to capture the increasing amplitude in the data.
@@ 3.2 Open-ended probabilistic model discovery for real-world datasets
Table 1: Comparison of LM programs against expert programs We perform this comparison across four different datasets and two different ablations that replace observations with synthetic observations and remove all metadata. We report the expected predictive log density estimated via leave-one-out cross validation. We bold statistically significant differences and underline non-significant differences. LM programs match the performance of human expert programs on 9/12 datasets.
By integrating LMs into the model discovery process, we can search over a much broader class of models. Here, we explore LMs ability to automatically construct probabilistic programs for datasets. We prompt a LM to produce probabilistic programs in pymc, a probabilistic programming language in Python, enabling us to search over a vast space of statistical models (Abril-Pla et al., 2023).
Dataset
We consider four real world datasets from the Stan PosteriorDB dataset (Magnusson et al., 2023): a dataset consisting of average improvement in SAT scores after a SAT improvement program across eight different high schools, a dataset consisting of ages of twenty-seven dugongs and their lengths, a surgical dataset consisting of mortality rates in twelve hospitals performing cardiac surgery on babies, and a dataset of peregrine population counts in the French Jura from 1964 to 2003. Each dataset has an associated human expert written probabilistic program in Stan, which we translate into pymc. These expert programs are generally open-source contributions from the Stan developer community. These datasets cover common modeling motifs such as hierarchical modeling and regression.
Ablations
To study how domain knowledge affects the LM’s modeling capabilities and to mitigate concerns about dataset leakage, we consider two ablations. In the first ablation, we construct a simulated analog for each dataset: for each dataset, we generate synthetic observations by sampling from the prior of a generative model or by sampling from a model with fixed parameters. We denote these datasets as simulated datasets. For example, to generate a simulated dataset for the eight schools dataset, we sample from the prior of the human expert program. To generate a simulated analog of the dugongs dataset, we fix the values of the parameters $\alpha,\beta,\gamma$ used in the expert model $y_{i}=\alpha-\beta\gamma^{x_{i}}$. Since it is highly unlikely that these simulated datasets appeared in GPT-4’s training data, this ablation helps mitigate concerns about dataset leakage.
In the second ablation, we remove dataset metadata from the information provided to the LM. In particular, we remove the dataset description, replace the column names with domain-agnostic column names (e.g., $x_{0},y$) and replace the axis labels with uninformative labels. We denote these datasets as no metadata datasets. By removing this metadata, we can characterize how LMs use domain knowledge and mitigate concerns about dataset leakage.
Quantitative Results
We compare the performance of our method across different datasets and ablations against expert programs in Table 1. We emphasize that the expert programs are strong baselines especially for the simulated datasets since the expert programs generated the data. We highlight numbers corresponding to significant differences, where significance is defined as an ELPD difference of greater than four times the standard error estimate.
LMs reliably identify programs on par with expert programs; we validate the fidelity of the posterior distributions using standard Markov Chain Monte Carlo diagnostics. Crucially, removing metadata or switching to simulated datasets does not generally reduce performance relative to the expert program. The main exception is the surgical dataset. By replacing the labels with $x_{0}$ and $y$, the LM formulates this problem as a regression problem instead of using a hierarchical model like the expert model. For the Peregrine dataset, removing metadata improves performance. We discuss this further in the next section.
Qualitative Analysis: Language Models are Domain Experts
Earlier, we asked whether LMs can reliably play the role of a domain expert. In our ablations, we study how metadata influences the model search process by examining how removing metadata changes the programs proposed by LMs. In Figure 3, we plot the model fit and list the corresponding LM programs in the figure caption. For the peregrine population dataset, when given all the metadata, the LM models this dataset using a logistic growth model, which is motivated by the problem domain. However, logistic growth models cannot capture the initial decline in population. Interestingly, when we remove all metadata, the LM performs better, because it uses a quartic polynomial to model the data. This closes the gap in performance with the expert program. We observe a similar trend with the dugongs growth curve dataset. When given metadata, the LM uses a von Bertalanffy growth function (von Bertalanffy, 1949), which is commonly used to model animal growth. When we remove all metadata, the LM uses a polynomial function with a logarithmically transformed value for the inputs.
Figure 3: Domain knowledge shapes LM modeling approaches. In the top row, we keep all metadata (e.g., dataset description). In the bottom row, we remove metadata that reveals information about the domain. This leads to qualitatively different approaches for three datasets; for eight schools, the LM discovers a hierarchical model even without metadata. We list the corresponding programs for these different ablations. (top row) $\frac{K}{(1+((K-P_{0})/P_{0})\exp(-rt))}$, $\text{BetaBin}(n,\alpha,\beta)$, $L_{\text{inf}}(1-\exp(-k(\text{age}-t_{0})))$; (bottom row) $a+bx_{0}+cx_{0}^{2}+dx_{0}^{3}+ex_{0}^{4}$, $\alpha+\beta x_{0}+\gamma x_{0}^{2}$, $\alpha+\beta_{1}\log x_{0}+\beta_{2}\log x_{0}^{2}+\beta_{3}\log x_{0}^{3}$.
Interestingly, the LM sets the prior parameters in these models in a data-informed way. For example, the LM sets the asymptotic length in the von Bertanlanffy function based on the observed lengths in the dataset. See code snippet in Figure 7 of the Appendix. For the eight schools dataset, the LM identifies a hierarchical model even without metadata, illustrating that model criticism can compensate for lack of domain knowledge. We illustrate the improvement round to round in Figure 4.
Figure 4: Round-to-round improvement. Model criticism leads to improvements on average, although improvement is not necessarily monotonic for a given dataset and run. (left) LOO score vs round for Stan dataset. We normalize the scores across a run of Box’s loop and average across Stan datasets for the no metadata, simulated condition. Larger is better and error bars correspond to standard error. (right) Squared error vs round for Lotka-Volterra experiment. We report the squared error averaged across three different runs, for the warm-start, no constraint condition; smaller is better.
@@ 3.3 Improving classic models under modeling constraints
Figure 5: Correcting misspecified Lotka-Volterra dynamics. LM can introduce corrections to standard Lotka-Volterra dynamics (no warm-start) and a hybrid neural approach (warm-start) that outperform several baselines. (left) LM proposed model predictions on training data and extrapolations (grey region); (right) LM models (warm-start and no-warm start) achieve lower test MAE than a neural ODE model, a hybrid ODE model with a multiplicative correction to the prey-predator dynamics, and the standard Lotka-Volterra (LV) model.
1def no_warm_start(t, y, _coeffs): 2  # Hypothesis 1: Logistic growth for prey with carrying capacity kappa 3  # Hypothesis 2: Saturation effect in predation rate with parameter psi 4  db_dt = alpha * b * (1 - b / kappa) - beta * b * c / (1 + psi * b) 5  dc_dt = -gamma * c + delta * b * c - epsilon * c**2 1def warm_start_constrained(t, y, _mlp): 2  inputs = jnp.array([b])  # Input is prey density (bunny population) 3  # Neural network learns a modifier for the handling time based on prey density 4  handling_time_modifier = _mlp(inputs) 5  # Modulate the predation rate beta in a non-linear manner according to handling time 6  db_dt = alpha * b - beta * b * c / (1 + handling_time_modifier[0] * b) 7  dc_dt = -gamma * c + delta * b * c / (1 + handling_time_modifier[0] * b)
Figure 6: LM can propose corrections to ODEs (top) In the no warm-start variation, the LM introduces corrections informed by domain knowledge of predator-prey models (carrying capacity, predation saturation) (bottom) When prompted to introduce neural networks in an interpretable way, one strategy the LM proposes is to make the handling time parameter depend non-linearly on the prey density, extending a traditional approach to modeling predation saturation.
In the previous experiment, we explored LM’s ability to identify models tabula rasa (e.g., without any initial seed model). However, in many scientific settings, we begin with a well-known model and are tasked with improving it. Here, we explore LM’s ability to improve upon a Lotka-Volterra model of predator-prey dynamics. We also explore LM’s ability to respect “soft” modeling constraints expressed in natural language (e.g., this model should be interpretable to ecologists), an important component of model discovery.
@@
$\frac{db}{dt}$ $=\alpha b-\beta bc$ (1)
$\frac{dc}{dt}$ $=-\gamma c+\delta bc^{0.95}$ (2)
@@
The LM is tasked with implementing a ODE model in Python using Jax (Bradbury et al., 2018). Estimating the parameters of Lotka-Volterra models via Bayesian inference is challenging. We instead learn the parameters via gradient descent which can be straightforwardly implemented using modern automatic differentiation libraries (Chen et al., 2018; Kidger, 2021).
We consider three variations: warm-start with constraints, warm-start with no constraints and no warm-start that differ in their initial seed program and the initial instructions which express modeling constraints. In all variations, we provide the LM with a scatter plot of training datapoints. In the no warm-start (WS) variation, we provide the LM with an implementation of standard Lotka-Volterra dynamics using diffrax (Kidger, 2021), a Jax based ODE library that supports learning ODE parameters via backpropagation and the predictions obtained from fitting standard Lotka-Volterra to the training data. In both warm-start (WS) variations, we provide the LM with an implementation of a hybrid Neural ODE that introduces an additive correction term to the predator dynamics, parameterized by a multilayer perceptron (MLP), and the predictions obtained from fitting this model to the training data. In the constrained warm-start variation, we ask the LM to produce a hybrid neural ODE model that is interpretable to an ecology expert. In the unconstrained warm-start variation, we provide the same seed program but we do not impose this additional constraint. Importantly, for models with both neural and physical components, we employ a two-stage learning procedure so that the neural component does not dominate the dynamics; see Appendix C for details.

### v2

@@ 3.1 Searching over a DSL: automated Gaussian process kernel discovery
Dataset BoxLM+ BoxLM Periodic AS SM N-BEATS
@@
Figure 2: Test set performance on time series datasets. Our BoxLM system identifies compositional kernels with performance on par with strong baselines. (left) Comparison of BoxLM test mean absolute error (MAE) against Automatic Statistician using greedy search (AS), spectral mixture kernel (SM), periodic kernel (Periodic), and N-BEATS. BoxLM+ searches over an augmented kernel space. We bold the best and underline the second best among the GP methods, treating N-BEATS as a powerful non-GP-constrained baseline. (right) Extrapolations from GP with a BoxLM-discovered kernel.
We first evaluate LM’s ability to search over a constrained space of models; in some settings, a domain expert may require the modeler to search over a DSL for reasons such as interpretability. In particular, we consider time-series modeling with Gaussian processes (GPs) as in the Automatic Statistician (Duvenaud et al., 2013). GPs are probabilistic models that specify a distribution over functions. GPs are defined by a mean function and a positive-definite kernel function $k(\mathbf{x},\ \mathbf{x^{\prime}})$ that gives the covariance between $f(\mathbf{x})$ and $f(\mathbf{x^{\prime}})$ as a function of $\mathbf{x}$ and $\mathbf{x^{\prime}}$. The properties of a GP depend significantly on the kernel and therefore choosing a kernel that reflects domain knowledge (e.g., linearity, periodicity) is a key design choice. One common way to produce a more flexible kernel is to compose kernels via addition and multiplication, leveraging the closure properties of kernels. Here, we evaluate LMs’ ability to search over a space of kernels to identify an appropriate composition of kernels.
@@
Mirroring Duvenaud et al. (2013), we define a space of base kernels. In the prompt, we ask $q_{LM}$ to perform one of the three operations (addition, multiplication, and replacement) on one of the in-context exemplar programs. For the first round, we ask $q_{LM}$ to produce an initial guess based on the structure of the dataset, which we provide as a plot. Given a kernel expression, we learn the kernel hyperparameters via gradient-based optimization of the marginal likelihood.
@@ Results
We evaluate our method on six common univariate time series datasets. We compare against the Automatic Statistician, a greedy algorithm proposed by Duvenaud et al. (2013) (run until a depth of 10), and two GP baselines: a periodic kernel and a spectral mixture kernel, a strong non-compositional baseline. We also compare against N-BEATS, a strong neural baseline (Wilson & Adams, 2013; Oreshkin et al., 2019). In Figure 2, we compare the mean absolute error (MAE) on held-out test data for all datasets. To account for stochasticity in the Automatic Statistician implementation we used (Saad et al., 2023) and stochasticity in different repetitions of our pipeline, we average the test MAE across three model runs. Our method, denoted LM in the table, matches the performance of the Automatic Statistician, showing that BoxLM can efficiently search over a constrained space of models. We also experiment with augmenting the base kernel space with additional kernels (denoted BoxLM+ in the table). In some cases, this additional flexibility is beneficial. For example, by using the additional kernels, BoxLM+ can much better capture the Australian beer sales dataset than LM; in other cases, the additional flexibility does not appreciably improve performance. On the right panel, we show the extrapolations and identified kernel for the monthly air passenger dataset; BoxLM+ identifies a kernel with a periodic times linear component to capture the increasing amplitude in the data.
@@ 3.2 Open-ended probabilistic model discovery for real-world datasets
By integrating LMs into the model discovery process, we can search over a much broader class of models. Here, we explore BoxLMs’ ability to automatically construct pymc probabilistic programs (Abril-Pla et al., 2023) for datasets.
Dataset
We consider four real world datasets from the Stan PosteriorDB dataset (Magnusson et al., 2023): (1) a dataset consisting of average improvements in SAT scores after an SAT improvement program across eight different high schools, (2) a dataset consisting of ages of twenty-seven dugongs and their lengths, (3) a surgical dataset consisting of mortality rates in twelve hospitals performing cardiac surgery on babies, and (4) a dataset of peregrine population counts in the French Jura from 1964 to 2003. Each dataset has an associated human expert written probabilistic program in Stan, which we translate into pymc. These expert programs are generally open-source contributions from the Stan developer community. The datasets cover common modeling motifs, such as hierarchical modeling and regression.
Figure 3: Domain knowledge shapes BoxLM modeling approaches. In the top row, we keep all metadata (e.g., dataset description). In the bottom row, we remove metadata that reveals information about the domain. This leads to qualitatively different approaches for three datasets; for eight schools, BoxLM discovers a hierarchical model even without metadata. We list the corresponding programs for these different ablations: (top row) $\frac{K}{(1+((K-P_{0})/P_{0})\exp(-rt))}$, $\text{BetaBin}(n,\alpha,\beta)$, $L_{\text{inf}}(1-\exp(-k(\text{age}-t_{0})))$; (bottom row) $a+bx_{0}+cx_{0}^{2}+dx_{0}^{3}+ex_{0}^{4}$, $\alpha+\beta x_{0}+\gamma x_{0}^{2}$, $\alpha+\beta_{1}\log x_{0}+\beta_{2}\log x_{0}^{2}+\beta_{3}\log x_{0}^{3}$.
Ablations
To study how domain knowledge affects the LM’s modeling capabilities and to mitigate concerns about dataset leakage, we consider two ablations. In the first ablation, we construct a simulated analog for each dataset: for each dataset, we generate synthetic observations by sampling from the prior of a generative model or by sampling from a model with fixed parameters. We denote these datasets as simulated datasets. For example, to generate a simulated dataset for the eight schools dataset, we sample from the prior of the human expert program. In some cases, such as the dugongs dataset, the prior distributions over parameters are highly unconstrained and we therefore fix values for the parameters instead. For example, to generate a simulated analog of the dugongs dataset, we fix the values of the parameters $\alpha,\beta,\gamma$ used in the expert model $y_{i}=\alpha-\beta\gamma^{x_{i}}+\epsilon_{i}$ where $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$. Since it is unlikely that these simulated datasets appeared in the training data, this ablation helps mitigate concerns about dataset leakage.
In the second ablation, we remove the dataset metadata from the LM prompt. In particular, we remove the dataset description, replace the column names with domain-agnostic column names (e.g., $x_{0},y$), and replace the axis labels with uninformative labels. We denote these datasets as no metadata datasets. By removing this metadata, we can characterize how LMs use domain knowledge.
Quantitative Results
In Table 1, we compare the performance of our method across different datasets and ablations relative to the expert programs. We emphasize that the expert programs are strong baselines, especially for the simulated datasets where the expert programs generated the data. We highlight numbers corresponding to significant differences, where significance is defined as an ELPD LOO difference of greater than four times the standard error estimate.
BoxLM reliably identifies programs on par with expert programs; we validate convergence using standard Markov Chain Monte Carlo diagnostics. Interestingly, removing metadata or switching to simulated datasets does not generally reduce performance relative to the expert program. The main exception is the surgical dataset. By replacing the labels with $x_{0}$ and $y$, BoxLM formulates this problem as a regression problem instead of using a hierarchical model like the expert model. For the peregrine dataset, removing metadata improves performance. We discuss this further in the next section.
Qualitative Analysis: Language Models are Domain Experts
Earlier, we asked whether LMs can reliably play the role of a domain expert. In our ablations, we study how metadata influences the model search process by examining how removing metadata changes the programs proposed by BoxLM. In Figure 3, we plot the model fit and list the corresponding BoxLM programs in the figure caption. For the peregrine dataset, when given all the metadata, BoxLM models this dataset using a logistic growth model, which is motivated by the problem domain. However, logistic growth models cannot capture the initial decline in population. Interestingly, when we remove all metadata, BoxLM performs better, because it uses a quartic polynomial to model the data. This closes the gap in performance with the expert program. This highlights how prior knowledge can have a (sometimes overly) strong influence on the modeling choices of BoxLM. We observe a similar trend with the dugongs growth curve dataset. When given metadata, BoxLM uses a von Bertalanffy growth function (von Bertalanffy, 1949), which is commonly used to model animal growth. When we remove all metadata, BoxLM uses a polynomial function with a logarithmically transformed value for the inputs. Here, both modeling approaches fit the data well. Interestingly, BoxLM sets the prior parameters in these models in a data-informed way. For example, BoxLM sets the asymptotic length in the von Bertanlanffy function based on the observed lengths in the dataset. See code snippet in Figure 7 of the Appendix. For the eight schools dataset, BoxLM identifies a hierarchical model even without metadata, illustrating that model criticism can compensate for lack of domain knowledge. We illustrate the improvement round to round in Figure 6.
@@
Table 1: Comparison of BoxLM programs against expert programs We perform this comparison across four different datasets and two different ablations that replace observations with synthetic observations and remove all metadata. We report the expected predictive log density estimated via leave-one-out cross validation. We bold statistically significant differences and underline non-significant differences. LM programs match the performance of human expert programs on 9/12 datasets.
@@ 3.3 Improving classic models under modeling constraints
Figure 4: Correcting misspecified Lotka-Volterra dynamics. BoxLM can introduce corrections to standard Lotka-Volterra dynamics (no warm-start) and a hybrid neural ODE approach (warm-start) that outperform several baselines. (left) LM-proposed model predictions on training data and extrapolations (grey region). (right) Test MAE of LM models (No-WS, WS-Constraint, and WS-No Constraint) compared to the standard Lotka-Volterra model LV, a Neural ODE, and a hybrid Neural ODE model with a multiplicative correction to the prey-predator dynamics (Hybrid).
LM proposed Lotka-Volterra programs 1def no_warm_start(t, y, _coeffs): 2 # Hypothesis 1: Logistic growth for prey with carrying capacity kappa 3 # Hypothesis 2: Saturation effect in predation rate with parameter psi 4 db_dt = alpha * b * (1 - b / kappa) - beta * b * c / (1 + psi * b) 5 dc_dt = -gamma * c + delta * b * c - epsilon * c**2 1def warm_start_constrained(t, y, _mlp): 2 inputs = jnp.array([b]) # Input is prey density (bunny population) 3 # Neural network learns a modifier for the handling time based on prey density 4 handling_time_modifier = _mlp(inputs) 5 # Modulate the predation rate beta in a non-linear manner according to handling time 6 db_dt = alpha * b - beta * b * c / (1 + handling_time_modifier[0] * b) 7 dc_dt = -gamma * c + delta * b * c / (1 + handling_time_modifier[0] * b) 1def warm_start_unconstrained(t, y, _mlp): 2 inputs = jfu.ravel_pytree((b, c))[0] 3 mlp_output = _mlp(inputs) 4 # Fine-tuned scaling of the MLP output to match the amplitude of data more closely 5 db_dt = alpha * b - beta * b * c + 0.02 * mlp_output[0] # Reduced scaling for bunnies 6 dc_dt = -gamma * c + delta * b * c + 0.06 * mlp_output[1] # Increased scaling for cats
Figure 5: BoxLM can propose corrections to ODEs. (top) In the no warm-start (No-WS) variation, BoxLM introduces corrections informed by domain knowledge of predator-prey models (carrying capacity, predation saturation). (middle) When prompted to introduce neural networks in an interpretable way (WS-Constraint), one strategy BoxLM proposes is to make the handling time parameter depend non-linearly on the prey density, extending a traditional approach to modeling predation saturation. (bottom) When prompted to introduce neural networks without constraints (WS-No Constraint), BoxLM introduces additive MLP-parameterized corrections and adjusts the scaling factors.
In the previous experiment, we explored BoxLM’s ability to identify models tabula rasa (e.g., without any initial seed model). However, in many scientific settings, we begin with a well-known model and are tasked with improving it. Here, we explore BoxLM’s ability to improve upon a Lotka-Volterra model of predator-prey dynamics. In addition, a crucial component of model discovery is respecting “soft” modeling constraints that are easy to express in natural language but hard to formalize (e.g., ecologists should think this is a plausible model). We therefore illustrate another advantage of BoxLM: we can express these modeling constraints in natural language and use them to drive LM proposals.
@@
$\frac{db}{dt}$ $=\alpha b-\beta bc$
$\frac{dc}{dt}$ $=-\gamma c+\delta bc^{0.95}.$
@@
BoxLM is tasked with implementing an ODE model in Python using Jax (Bradbury et al., 2018). Estimating the parameters of Lotka-Volterra models via Bayesian inference is challenging. We instead learn the parameters via gradient descent which can be straightforwardly implemented using modern automatic differentiation libraries (Chen et al., 2018; Kidger, 2021); in particular, we use diffrax (Kidger, 2021), a Jax-based ODE library that supports learning ODE parameters via backpropagation. We consider three variations: warm-start with constraints (WS-Constraint), warm-start with no constraints (WS-No Constraint), and no warm-start (No-WS) that differ in their initial seed program and the initial instructions which express modeling constraints. In all variations, we provide the LM with a scatter plot of training datapoints. In the no warm-start variation, we provide the LM with an implementation of standard Lotka-Volterra dynamics using diffrax and the predictions obtained from fitting standard Lotka-Volterra to the training data. In both warm-start variations, we provide the LM with (1) an implementation of a hybrid neural ODE that introduces an additive correction term to the predator dynamics, parameterized by a multilayer perceptron (MLP), and (2) the predictions obtained from fitting this model to the training data. In the constrained warm-start variation, we ask the LM to produce a hybrid neural ODE model that is interpretable to an ecology expert who suggested a Holling’s type II response (Rosenzweig & MacArthur, 1963). In the unconstrained warm-start variation, we provide the same seed program but do not impose this additional interpretability constraint. For models with both neural and physical components, we employ a two-stage learning procedure so that the neural component does not dominate the dynamics; see Appendix C for details.